# PDMS elastic properties: influence of fabrication protocol and test method

### v1

@@ 2.3 Correlation analysis
Compiled data is visualized to observe trends and the monotonicity of relationships is quantified using Spearman rank correlation coefficients, $\rho$ [62]. The Spearman correlation is executed via MATLAB and is applied between preparation conditions (cure temperature, time) and within subgroups of individual test methods in order to isolate possible origins of stiffness discrepancies. While $\rho$ provides indication of how monotonic the relation between two inputs is ($0.60.3\leq|\rho|<$ 0.6: Fair, $0.80.6\leq|\rho|<$ 0.8: Moderate, $10.8\leq|\rho|<$ 1: Very strong, $|\rho|=1$: Perfect[63]), it does indicate the magnitude or linearity of an effect.
Figure 2: Stiffness data from all included studies relative to material composition show wide variability in reported moduli. Sorting by test methodology reveals trends in how test methods correspond to mix ratios. The stacked histogram shows that the TDS recommended mix ratio of 10:1 wtbase:wtxlinker has the most reported data for moduli in literature. The abundance of data at 10:1 makes trend interpretation for combined data challenging when studies reporting only a value are averaged in with studies exploring stiffness by mix ratio. Tension testing is common for stiffer samples however cavity expansion dominates in softer mix ratios while Nanoindentation and compression have been used across a range of compositions.
3 Results
@@ 3.1 Wide modulus variability and dominance of 10:1 ratio
Despite reportedly identical mix ratios, the measured shear moduli can vary by three orders of magnitude between different studies as shown in Fig. 2. We observe that this is due to three main factors: 1) different test methods producing ranging results, 2) variation in specified cure schedule, and 3) the inclusion of non reactive oils in the base mixture that are sometimes included to manipulate the viscoelasticity of the cured silicone. The histogram in Fig. 2 shows a count of how many experimental results exist by test method across a range of crosslinking densities and highlights how different test methods tend to be used for different stiffness of materials. While combining all the data is useful to inspire caution when using literature reported “stiffness” for SYLGARD 184, diving further into the data is necessary to try and identify the sources of the variability and the key parameters driving the reported ranges. Note that in the Technical Data Sheet (TDS), Dow reports a tensile modulus of 6.7 MPa for a 10:1 mix ratio[15]. Though not included in Fig. 2 given that no test method or cure conditions are available, this is significantly above all results compiled here.
Table 1 presents the Spearman correlation between shear modulus and key process parameters: curing agent ratio, cure temperature, and cure time. Correlations are calculated for all tests and cure ratios combined, then separately for select testing methods with data from multiple authors (and conditions). The second section of the table is split into only tests performed at the 10:1 ratio (that recommended on the product TDS [15]), and those excluding this ratio. All comparisons between stiffness and curing agent ratio indicate very strong correlations with an exceptionally low p-value, as expected, except for those tests performed using tension testing, which returns only fair to moderate correlation.
3.2 Hotter cures result in stiffer material
Table 1 also shows the counterintuitive result that longer cure times correlate with softer samples, however this is likely due to the significantly longer times used with room temperature curing. With the strength of the correlation between temperature and stiffness, this can be explained by the tendency of lower temperature curing to produce softer samples overshadowing the effect of time on cure. In an effort to reduce the variability imparted by different test methods, Table 1 is subdivided by test types and further to isolate 10:1 ratio samples. Across compression, nanoindentation and tension testing, the correlation between cure temperature and stiffness increases to moderate - strong when test methods are considered independently, and is even stronger when 10:1 ratio samples are excluded (confirming that the number of tests performed only at this ratio may be biasing the overall trends in results).
There is, however, a limit at which higher temperatures will no longer produce stiffer samples and instead will degrade the polymer. Liu et al. [39] set out to study the effect of heating on PDMS mechanical properties and also observe that stiffness increase between 100 and 150 C, but decreases for 200 C and 300 C samples. [^2] The data displayed here from [39] have been down selected to only the 100 C and 200 C data, though 150 C and 300 C cure temperatures were also evaluated in the original work.
@@ 3.2 Hotter cures result in stiffer material
Table 1: Correlations of stiffness with preparation conditions ($\rho$) and p-value of each. Very strong negative correlations between curing agent ratio and modulus across all test methods and groups of samples appears as expected. The later two sections of the table isolate samples prepared at the 10:1 w/w in order to remove the disproportionate number of tests recorded at this ratio from biasing the results. Blanks in the table indicate cases where conditions are the same for all trials within the row.
Figure 3: Cure temperatures are not equally likely across all wB+oil:wC ratios. Room temperature cures are more common for larger ratio samples and cure temperatures over 150 C are rarely seen for mix ratios over 10:1.
@@ 3.3 False equivalencies arise due to test method
In an effort to isolate the cure schedule from the test method, Fig. 4 compares stiffness for samples cured 120 minutes at 85-100 C across multiple authors. Fitting this data with an exponential function results in $\mu=1003e^{-0.073w_{B+oil}}$ with an R2 of 0.74 (a linear fit on the same region had an R2 of 0.65). Particularly for the higher curing agent ratio samples, the measured stiffness varies by hundreds of kilopascals between testing methods. Tension testing inconsistently reports moduli both higher and lower than nanoindentation, and is not used for softer samples. While the general softening trend with higher base ratios can provide a rough qualitative understanding of polymer composition’s influence on stiffness, there is minimal overlap in test methods that are used between soft samples and stiff samples, leaving open the question that the trends are due to the test methodology.
@@
Figure 4: Test method imparts variability even when a similar cure schedule is used: 120 minutes at between 85 and 100C. Test type indicated by marker shape matching those used in Figure 2.
@@
Restricting the wB:wC ratio to a more narrow range and examining three cavity-expansion based methods, encourages caution about biases that may be more pervasive in the dataset. Fig. 5 compares the $\mu_{0}$ of Yang et al. [57] measured through CR and assuming a neo-Hookean material model, to results obtained via VCCE for comparable mix ratios (oil containing samples are excluded). The VCCE results from Raayai-Ardakani et al. [17, 52] are also $\mu_{0}$, while those of Chockalingam et al. [53] and this work are $\mu$.
@@
As expected, $\mu<\mu_{0}$ when measured with VCCE with instantaneous moduli more than double the quasistatic moduli for comparable wB:wC ratio. However, it can be seen on Figure 5 that $\mu_{0}$ does not correlate 1:1 between CR and VCCE. The observed deviation between Raayai-Ardakani and Yang et al. can therefore be attributed to the 6 day room temperature cure of Yang et al. producing softer samples than the 3 days at 40 C cure schedule used by Raayai-Ardakani. By contrast, the unexpected 1:1 correlation between $\mu_{0}$ by CR [57] and $\mu$ by VCCE [53] (blue line on Figure 5) is likely attributed to the discrepancy between the room temperature cure schedule used byt Yang et al. and the 100 C 2 hr schedule used by Chocaklingam et al. and this work.
@@
Figure 5: VCCE results of [17, 52] for $\mu_{0}$ (filled marker), as well as $\mu$ from [53] and this work (open circle, blue axis labels) compared to CR results for $\mu_{0}$ by Yang et al. [57] for comparable mix ratios (wB:wC as noted on plot).
3.4 Nonreactive oils act as part of the $w_{B}$ in stiffness trends
@@
Oil-based thinners have been proposed as a method to decouple the viscoelastic and elastic responses of PDMS without effecting the stiffness. An unspecified ‘thinner’ was used in [36] while [44, 53, 57], and new data presented in this work, use silicone oil added to the base. Fig. 6 again shows results obtained by Yang et al. alongside those of Chockalingam et al. and this work, demonstrating that the correlation of shear modulus to mix compositions is stronger when the added oil is included in the $w_{B}$ than when it is not ($|\rho|=0.96$ vs $|\rho|=0.47$ in [57] and $|\rho|=0.85$ vs $|\rho|=0.74$ combining [53] and this work). Across both sets of data, increasing $w_{B+oil}:w_{C}$ reduces the observed shear modulus ($\mu$ or $\mu_{0}$) by a similar amount (0.768 vs 0.766 kPa per added $w_{B+oil}$ with a linear fit to the data in this region). While all authors report changes to the viscoelastic time constants when including nonreactive oils, the trend in softening with increasing $w_{B+oil}$ for soft PDMS is shown to be consistent a) regardless as to whether the ratio is altered with base or oil, and b) in both instantaneous and quasistatic modulus.
@@
Figure 6: If wB:wC is considered without taking oil into account (crosses), the correlation of ratio with stiffness is not as strong as when wB+oil:wC is considered (circles) and the R2 of a linear fit drops from 0.88 to 0.21 for CR and 0.83 to 0.31 for VCCE. The top figure shows $\mu_{0}$ from Yang et al. [57] against an abscissa calculated both with and without oil included in $w_{B}$; the bottom plot show $\mu$ results from Chockalingam et al., [53] and this work.
@@ 4 Conclusions
It is clear that current literature is far from providing reproducible scientific results when it comes to measuring the mechanical properties of SYLGARD 184. If this is our goal as a community, a number of steps must be taken to better observe the properties that so many authors to date have claimed to measure. At present, cure schedules ranged widely between researchers with very little overlap if the authors were not currently or historically from the same research group. Of the utmost importance, is unifying how PDMS is cured and reporting full details of these conditions, as well as timing between cure and test. We would propose three temperatures for PDMS where control of mechanical properties is desired: a $25^{\circ}$ room temperature, a “moderate” $60^{\circ}$ at or below the glass transition temperature of plastics commonly used in additive manufacturing, and a “hot” $100^{\circ}$ for rapid curing.
While the TDS recommends four possible cure schedules [15] between room temperature and 150 C, numerous authors note changes in mechanical properties after “fully curing” material with the recommended conditions. Further attention on storage time and condition is warranted as pointed out by a number of the studies including [53] and [48] who showed nearly a doubling in stiffness with 18 vs 3 days of storage at room temperature after an initial cure.
@@
Tension tests on PDMS reported the widest variation of stiffness, with nanoidentation also providing widely variable results. Though these are both common testing techniques, we recommend caution in the use of both methods. At low mixing ratios PDMS is brittle [42] and tension testing is highly sensitive to surface imperfections on the sample, while at higher wB:wC ratios, the soft modulus means measurement set up is prohibitively challenging and results are influenced by the clamping conditions. The variability of nanoindentation is likely due to the uncertain nature of the adhesion between the sample and indentor tip, which was the focus of many of the authors referenced here. Using a method that relies on consistent surface adhesion, in a material where surface treatments are applied to widely varying degrees, will continue to produce imprecise results.
@@
Each researcher has unique needs for their research, and not each mechanical testing method is possible for a given sample. However, by noting the current discrepancies in literature and advocating additional attention is focused on consistency and comparisons to the full body of literature going forward (not selected works that happen to align with ones results), we anticipate better agreement in measured mechanical properties is easily attainable if the community decides it is important.

### v2

@@ 2.3 Correlation analysis
Compiled data is visualized to observe trends and the monotonicity of relationships is quantified using Spearman rank74 correlation coefficient, $\rho$. The Spearman correlation presents the null hypothesis that there is no correlation between an input (predictor) variable and the output, and returns $-1<\rho<1$ indicating whether a monotonic correlation exists, and whether it is negative or positive, respectively. A significance p$<0.05$ indicates confidence in rejecting the null hypothesis of the data originating from an uncorrelated dataset (more precisely, from a dataset that is less well correlated than indicated by the magnitude of $\rho$). Spearman correlation is used given the non-normal distribution of all predictor variables, and the desire to not impose a linear relationship between the inputs and stiffness 75. Further discussion of the statistical distribution of the predictor variables is given in Sec. 4 of the ESI.
The correlation is executed via MATLAB 2023 and is applied between preparation conditions (cure temperature, time) and within subgroups of individual test methods in order to isolate where monotonicity might hold internal to a test method, but not within the entire population. While $\rho$ provides indication of how monotonic the relation between two inputs is ($0.60.3\leq|\rho|<$ 0.6: Fair, $0.80.6\leq|\rho|<$ 0.8: Moderate, $10.8\leq|\rho|<$ 1: Very strong, $|\rho|=1$: Perfect76), it does not indicate the magnitude or linearity of an effect.
Fig.  2: Stiffness data from all included studies relative to material composition show wide variability in reported moduli. Sorting by test methodology reveals trends in how test methods correspond to mix ratios. The stacked histogram shows that the TDS recommended mix ratio of 10:1 wtbase:wtxlinker has the most reported data for moduli in literature. The abundance of data at 10:1 makes trend interpretation for combined data challenging when studies reporting only a value are averaged in with studies exploring stiffness by mix ratio. Tension testing is common for stiffer samples, however Cavity Expansion dominates in softer mix ratios while Nanoindentation and Compression have been used across a range of compositions.
3 Results & Discussion
@@ 3.1 Wide modulus variability and dominance of 10:1 ratio
Despite reportedly identical mix ratios, the measured shear moduli vary by two orders of magnitude between different studies as shown in Fig. 2. We observe that this is due to three main factors: 1) different test methods producing ranging results, 2) variation in specified cure schedule, and 3) the inclusion of non-reactive oils in the base mixture that are sometimes added to manipulate the viscoelasticity of the cured silicone. The histogram in Fig. 2 shows a count of how many experimental results exist by test method across a range of curing agent ratios and highlights how different test methods tend to be used for different stiffness of materials. While combining all the data is useful to inspire caution when using literature reported “stiffness” for SYLGARD 184, diving further into the data is necessary to try and identify the sources of the variability and the key parameters driving the reported ranges.
[…]
Tension testing results tend to be less correlated than other test methods. All comparisons between stiffness and curing agent ratio indicate very strong correlations with an exceptionally low p-value, as expected from the reaction chemistry described in ESI.5. However, for those tests performed using tension testing, the correlation is only fair to moderate. Examining cure temperature, when all curing agent ratios are considered, the correlation with cure temperature is higher within tension testing than across other methods. However, at the 10:1 ratio where one would expect the dependency to be most cleanly expressed given the elimination of one variable entirely, this trend does not hold and the data fail the correlation with p=0.12. Cure time is weakly correlated when examining all tests combined. However, within individual test methods correlation routinely fails the null hypothesis and predicts a weaker or inverted correlation when evaluated with tension testing. Across all test methods and within tension testing, additional data are required to conclude that there is monotonic correlation.
Note that in the TDS, Dow reports a tensile modulus of 6.7 MPa for a 10:1 mix ratio35. Though not included in Fig. 2 or Table 1 given incomplete information on test method and cure conditions, this is significantly above all results compiled here.
Table 1: Correlations of stiffness with preparation conditions ($\rho$) and p-value of each. Very strong negative correlations between curing agent ratio and modulus across all test methods and groups of samples appears as expected. The later two sections of the table isolate samples prepared at the 10:1 wB:wC in order to remove the disproportionate number of tests recorded at this ratio from biasing the results. Blanks in the table indicate cases where conditions are the same for all trials within the row.
@@ 3.2 False equivalencies arise due to test method
Mix ratio and cure temperature are expected to consistently effect stiffness regardless of the test method. Indeed, examining Table 1, correlations of stiffness with curing agent ratio and temperature are observed within a given test method. However, when all methods are combined this correlation is weaker. This provides evidence of test method bias.
We hypothesize that the variation imparted by test method is driven by a number of factors including, but not limited to: material rate dependence, such that different methods measure the instantaneous vs relaxed modulus, the additional models required to separate stiffness from adhesion effects in contact measurements, and sample preparation and fixturing imparting imperfections in boundary conditions and defect nucleation.
@@
In an effort to isolate the influence of cure schedule from that of the test method, Fig. 3 compares stiffness for samples cured 120 minutes at 85-100 C across multiple authors. Fitting this data with an exponential function results in $\mu=1003e^{-0.073w_{B+oil}}$ with an R2 of 0.74 (a linear fit on the same region had an R2 of 0.65). Particularly for the higher curing agent ratio samples, the measured stiffness varies by hundreds of kilopascals between testing methods. Tension testing inconsistently reports moduli both higher and lower than nanoindentation, and is not used for softer samples. While the general softening trend with higher base ratios can provide a rough qualitative understanding of polymer composition’s influence on stiffness, there is minimal overlap in test methods that are used between soft samples and stiff samples, leaving open the question that the trends are due to the test methodology.
@@
Fig.  3: Test method imparts variability even when a similar cure schedule is used: 120 minutes at between 85 and 100C. Test type indicated by marker shape matching those used in Fig. 2.
@@
Restricting the wB:wC ratio to a more narrow range and examining three cavity-expansion based methods, encourages caution about biases that may be more pervasive in the dataset. Fig. 4 compares the $\mu_{0}$ of Yang et al. 69 measured through CR and assuming a neo-Hookean material model, to results obtained via VCCE for comparable mix ratios (oil containing samples are excluded). The VCCE results from Raayai-Ardakani et al. 11, 64 are also $\mu_{0}$, while those of Chockalingam et al. 65 and this work are $\mu$.
As expected from a solid mechanics perspective, $\mu<\mu_{0}$ when measured with VCCE with instantaneous moduli more than double the quasistatic moduli for comparable wB:wC ratio. However, in these datasets, all samples used to measure $\mu$ were cured hotter and for less time than all samples to measure $\mu_{0}$. This results in the inversion of $\rho$ in Table 1 for both cure temperature and time when considering VCCE independently. This inversion is another indication that the comparison of measured stiffness between different test methods and/or cure protocols can be misleading.
@@
Furthermore, it can be seen on Fig. 4 that $\mu_{0}$ does not correlate 1:1 between CR and VCCE. The observed deviation between Raayai-Ardakani and Yang et al. can therefore be attributed to the 6 day room temperature cure of Yang et al. producing softer samples than the 3 days at 40 C cure schedule used by Raayai-Ardakani. By contrast, the unexpected 1:1 correlation between $\mu_{0}$ by CR and $\mu$ by VCCE (blue line on Fig. 4) is likely attributed to the discrepancy between the room temperature cure schedule used by Yang et al. 69 and the 100 C 2 hr schedule used by Chocaklingam et al. 65 and this work.
@@
Fig.  4: VCCE results of 11, 64 for $\mu_{0}$ (filled marker), as well as $\mu$ from Chockalingam65 and this work (unfilled cyan circle, cyan axis labels) compared to CR results for $\mu_{0}$ by Yang et al. 69. Horizontal green lines specify the measured $\mu_{0}$ from cavitation rheology for comparable mix ratios (wB:wC as noted on plot)measured with VCCE. Dashed lines are visual aids depicting both a 1:1 correlation for $\mu$:$\mu_{0}$ and lack of correlation for $\mu_{0}$:$\mu_{0}$.
3.3 Hotter cures result in stiffer material
Fig.  5: Cure temperatures are not equally likely across all wB+oil:wC ratios. Room temperature cures are more common for larger ratio samples and cure temperatures over 150 C are rarely seen for mix ratios over 10:1.
Table 1 also shows the counterintuitive result that longer cure times correlate with softer samples, however this is likely due to the significantly longer times used with room temperature curing. While stiffness does not follow a linear trend with all predictor variables (as shown in Fig. 3) a basic regression performed on the complete data set provides insights into the relative importance of temperature, time, and curing agent ratio. Utilizing MATLAB 2023b to fit a linear regression model, cure temperature is connected to stiffness as 2.71 kPa/C (standard error SE of 0.74 kPa/C) while cure time is effected as 0.004 kPa/min cure (SE 0.006 kPa/min and a p$>$0.05, thus failing the assumption of a linear correlation). Curing agent ratio is most strongly linked to stiffness as -15.5 kPa/1 wB+oil (SE 1.7 kPa/1 wB+oil). Contour plots depicting the interplay between all three predictors can be seen in Sec. 6 of the ESI. With the strength of the correlation between temperature and stiffness, the tendency of lower temperature curing to produce softer samples overshadows the effect of longer cure times producing stiffer samples.
In an effort to reduce the variability imparted by different test methods, Table 1 is subdivided by test types and further to isolate 10:1 ratio samples. Across nanoindentation and tension testing, the correlation between cure temperature and stiffness increases to moderate - strong when test methods are considered independently, and is even stronger when 10:1 ratio samples are excluded (confirming that the number of tests performed only at this ratio may be biasing the overall trends in results).
There is, however, a limit at which higher temperatures will no longer produce stiffer samples and instead will degrade the polymer. Liu et al. 51 set out to study the effect of heating on PDMS mechanical properties and also observe that stiffness increase between 100 and 150 C, but decreases for 200 C and 300 C samples.The data displayed here from Liu at al. 51 have been down selected to only the 100 C and 200 C data, though 150 C and 300 C cure temperatures were also evaluated in the original work.
3.4 Nonreactive oils act as part of the wB in stiffness trends
Oil-based thinners have been proposed as a method to decouple the viscoelastic and elastic responses of PDMS without effecting the stiffness 69, 56. An unspecified “thinner” was used by Schneider et al. 46 while Moučka et al. 56 use 194 cSt silicone oil and Chockalingam et al. 65, Yang et al.  69 and the new results presented in this work use 350 cSt silicone oil added to the base.
@@ 3.4 Nonreactive oils act as part of the wB in stiffness trends
Fig. 6 shows that the correlation of shear modulus to mix compositions is stronger when the added oil is included in the $w_{B}$ than when it is not. The correlation for Yang et al. increased to $|\rho|=0.96$ from $|\rho|=0.47$ and $|\rho|=0.85$ from $|\rho|=0.74$ combining Chockalingam et al. and this work. Across both sets of data, increasing $w_{B+oil}:w_{C}$ reduces the observed shear modulus ($\mu$ or $\mu_{0}$) by a similar amount (0.768 vs 0.766 kPa per added $w_{B+oil}$ with a linear fit to the data in this region). While all authors report changes to the viscoelastic time constants when including nonreactive oils, the trend in softening with increasing $w_{B+oil}$ for soft PDMS is shown to be consistent a) regardless as to whether the ratio is altered with base or oil, and b) in both instantaneous and quasistatic modulus.
@@
Fig.  6: If wB:wC is considered without taking oil into account (crosses), the correlation of ratio with stiffness is not as strong as when wB+oil:wC is considered (circles) and the R2 of a linear fit drops from 0.88 to 0.21 for CR and 0.83 to 0.31 for VCCE. The top figure shows $\mu_{0}$ from Yang et al. 69 against an abscissa calculated both with and without oil included in $w_{B}$; the bottom plot show $\mu$ results from Chockalingam et al., 65 and this work.
3.5 Reaction chemistry offers hypotheses for variation
A number of the studies point out changes in mechanical properties of SYLGARD 184 within days to weeks after completing cure 65, 60. Cheng et al. showed nearly a doubling in stiffness with 18 vs 3 days of storage at room temperature after an initial cure. In their work with pure PDMS, Esteves et al. propose that secondary reactions occur as PDMS polymerizes 77. They observed that consumption of Si-H groups in the cross linker cannot be fully explained by hydrosylation (the conventionally understood pathway of PDMS cross linking) thus requiring that additional pathways are present. These additional pathways would be effected by ambient oxygen and moisture, pointing to relative humidity during fabrication as an additional parameter to consider in fabrication and storage of PDMS.
Recent work by Bardelli et al. 78 points to cure temperature as a factor in changing the molecular weight between crosslinks $M_{c}$ in PDMS. As described in ESI.5, $M_{c}$ is directly connected to the stiffness of a polymer. Bardelli et al. observe nearly a doubling of $\mu$ between a 4 hour cure at 65 C compared to 150 C (with $M_{c}$ decreasing from 2050 to 1050 g/mol). Additionally, if lower temperature cures allow for more secondary reactions to take place, this could explain the resulting solid being softer.
@@ 4 Conclusions
The choice of mechanical testing method can have a significant influence of the measured stiffness of soft materials. Using commercially available PDMS (SYLGARD 184) as a test case, we examine this bias by combining new experimental data and that surveyed from 32 published papers, spanning 9 different testing methods. Statistical analysis reveals that bias imparted by test method is compounded by inconsistent curing approaches. The current methods of both testing and reporting are thus far from providing reproducible results for the mechanical properties of SYLGARD 184 and, by extension, for other soft and biological materials. A number of steps must be taken to better observe the properties that so many authors to date have claimed to measure.
@@
Test method has been implicated previously as returning variable measurements within identical soft material samples, and PDMS is no exception. Tension tests on PDMS reported the widest variation of stiffness, with nanoindentation also providing widely variable results. Though these are both common testing techniques, we recommend caution in the use of both methods. At low mixing ratios PDMS is brittle 54 and tension testing is highly sensitive to surface imperfections on the sample, while at higher wB:wC ratios, the soft modulus means that results are influenced by the clamping conditions, without a guarantee that it is possible to set up a successful test at all. The variability of nanoindentation is likely due to the uncertain nature of the adhesion between the sample and indenter tip, which was the focus of many of the authors referenced here. Using such a method that relies on consistent surface adhesion, in a material where surface treatments are applied to widely varying degrees, will continue to produce imprecise results.
The limitations of testing make it challenging to characterize the composition-property relationship of SYLGARD 184. This difficulty is further accentuated by the inconsistency of fabrication conditions and reporting of curing protocols. At present, cure schedules ranged widely between researchers with very little overlap if the authors were not currently or historically from the same research group. Of utmost importance is fully reporting the cure schedule and preparation method, as well as timing between cure and test. Unifying how PDMS is cured to obtain well characterized responses could help to enhance data comparisons and reliability. Hence, we would propose three temperatures for PDMS where control of mechanical properties is desired: a $25^{\circ}$ room temperature, a “moderate” $60^{\circ}$ at or below the glass transition temperature of plastics commonly used in additive manufacturing, and a “hot” $100^{\circ}$ for rapid curing. Curing times may vary depending on sample geometry.
@@
Each researcher has unique needs and capabilities for their research, and not each mechanical testing method is possible for a given sample. As soft materials research continues to expand into more complex and biological systems that present additional challenges for reproducible measurements, building agreement around common synthetic soft materials is worthwhile for the community. By noting the current discrepancies in literature on PDMS and by advocating that work going forward place a greater emphasis on consistency and comparison to the full body of literature (not selected works that happen to align with ones results), we anticipate better agreement in measured mechanical properties is easily attainable and hope that this manuscript motivates this as a community goal.
Author Contributions
Hannah Varner: Writing - Original Draft, Writing - Review & Editing, Conceptualization, Investigation, Data Curation, Methodology, Formal Analysis, Visualization. Tal Cohen: Conceptualization, Writing - Review & Editing, Resources, Supervision.
Conflicts of interest
There are no conflicts to declare.